The Wall Street Journal reported on January 26, 2010, that Ford Motor Co. was going to announce the addition of a second shift at its Chicago plant that would create 1,200 jobs. However, under a new union agreement that allows a two-tiered wage system, as many as 600 of these workers will earn an average of $14/hour, which is about half of the current new hire wage.

  • What effect will this new lower wage rate have on Ford’s contribution margin?
  • How will it affect Ford’s break-even point?
Answer #1

New lower wage rate will increase the contribution margin.

This is because there is decrease in variable expense. So contribution margin increases.

Break even point will decrease.

If contribition margin increases , break even point will decrease.

For example

If fixed cost is 10000

Contribution margin is 10 , then break even point is = 10000/ 10 = 1000

If conttobution margin increases to 20 then

Break even point will = 10000/ 20 = 500

Break even point reduces.